OpenStudy (anonymous):

When a complex number is written in its polar form, z = r (cos theta + i sin theta), the angle theta is called the "argument" of _________.

OpenStudy (anonymous):

i just came here to say that I HATE complex numbers haha.. just because I don't know about the complex world.. except for i xD i^2 = -1 =)

OpenStudy (kropot72):

\[\theta\] is called the argument of z where z = a + ib